1. What is the budgeted sales price per unit? 2. What is the budgeted variable expense per unit? 3. What is the budgeted fixeThe Great Balloon Company Actual vs. Budget Performance Report For the Month Ended January 31 Master Master Budget Actual BudRequirement 1. What is the budgeted sales price per unit? The budgeted sales price per unit is $ |
Answer #1
Actual Master Budget Master Budget Variance
Sales Volume 57,000 54,000 3,000 F
Sales Revenue $     201,200 $        178,200 $        23,000 F
Less Variable Expenses $       75,300 $          64,800 $        10,500 U
Contribution Margin $     125,900 $        113,400 $        12,500 F
Less Fixed Expenses $       63,900 $          63,000 $              900 U
Operating Income $       62,000 $          50,400 $        11,600 F

Budgeted Sales price per unit = $178200 / 54000 = $3.30 per unit

Budgeted variable expense per unit = $64800 / 54000 = $1.20 per unit

Budgeted Fixed Cost per unit = $63000 / 54000 = $1.17 per unit